OpenStudy (dm91):

What is the value of the variable that makes the statement true? b + (–2) + 8 = 0 A. –10 B. –6 C. 6 D. 10

OpenStudy (mathstudent55):

First, add -2 and 8. What do you get?

OpenStudy (anonymous):

you want the left side to equal 0, so if 8 -2 = 6, what plus 6 would equal 0?

OpenStudy (dm91):

-10

OpenStudy (mathstudent55):

No. -2 is a negative number. -2 + 8 = 6 b + (–2) + 8 = 0 b + 6 = 0 Now subtract 6 from both sides.

OpenStudy (dm91):

thank you my brain is not working yet